OpenStudy (anonymous):

solve formula for the given letter y=3x+8 solve for x

OpenStudy (anonymous):

(y-8)/3=x. Just revers order of operations to move everything off of x and onto y.

OpenStudy (anonymous):

can you explain more

OpenStudy (anonymous):

Yeah, sure. So you know the order of operations right? PEDMAS? To remove numbers from one side of an equation, all you need to do is go in reverse of that. In other words, SAMDEP: y=3x+8. (SA) Subtract 8 from both sides y-8=3x. (MD) Now divide out the 3 (There is no E, but you add Parenthesis to show that the subtraction step is done first). (y-8)/3=x. Viola!